OpenStudy (anonymous):

Find the sum of the first 90 terms of the sequence –4, –1, 2, 5, 8

OpenStudy (anonymous):

–4, –1, 2, 5, 8 Here first term t1 = -4 common difference d= -1-(-4) =-1+4= 3 and n= 90 sum of nth term of an AP is given as: \[\huge S_n = \frac{n}{2} [2a+(n-1)d]\] Substituting the above values \[\huge S_{90} = \frac{90}{2} [2(-4)+(90-1)3]\] \[\huge \rightarrow S_{90} = 45 [-8+89 \times 3]\] \[\huge \rightarrow S_{90} = 45 [-8+267] = 45 \times 259\] \[\huge \rightarrow S_{90} =11655\] @AlexaNicole504

OpenStudy (anonymous):

here t1 =a